what is the slope of the line containing the points (3.4) and -6,10)
Answer by Maths68(1474)

Points (3,4) and (-6,10)
Slope (m) = (y2-y1)/(x2 - x1)
Slope (m) = (10-4)/(-6 - 3)
Slope (m) = 6/-9
Slope (m) = 2/-3
Slope (m) = - 2/3